OpenStudy (anonymous):

convert degrees Fahrenheit (F) into degrees Celsius (C) use the formula C = 5/9(F - 32) Rewrite the equation to convert degrees Celsius to degrees Fahrenheit.

OpenStudy (anonymous):

@JA1

OpenStudy (ja1):

Ok so we have the formula already: \[C = \frac{ 5 }{ 9 } (f - 32)\]

OpenStudy (anonymous):

A. F = 5/9(C - 32) B. F = 9/5 C - 32 C. F = 5/9 C - 32 D. F = 9/5 C + 32

OpenStudy (ja1):

So how can we make this do the opposite and convert C to F?

OpenStudy (anonymous):

Add, or multiply? By doing opposite.

OpenStudy (ja1):

Not quite, we simply have to switch the terms to be the opposite.

OpenStudy (ja1):

So first let's switch it term by term, the first term is: \[C =\] so we have to switch it to:\[F =\] since we are looking for F.

OpenStudy (anonymous):

C = 5/9 & F = - 32

OpenStudy (ja1):

Now the next term is: \[\frac{ 5 }{ 9 }\] and since we are doing the opposite we switch the fraction so what would the end product be?

OpenStudy (ja1):

Not quite.

OpenStudy (anonymous):

See, I am lost. (/.\)

OpenStudy (ja1):

xD ok so we just switched \(C~=\) into : \(F~=\) So what do we switch \[\frac{ 5 }{ 9 }\] into?

OpenStudy (anonymous):

F?

OpenStudy (ja1):

Remember we are doing the opposite.

OpenStudy (ja1):

Not quite, just reorder the fraction to be the opposite :)

OpenStudy (ja1):

5/9 opposite = ?

OpenStudy (anonymous):

9/5

OpenStudy (ja1):

Exaclty :)

OpenStudy (anonymous):

So, it's B.

OpenStudy (ja1):

So we now we have: \[F = \frac{ 9 }{ 5 }\]

OpenStudy (ja1):

Not yet...

OpenStudy (anonymous):

no, its D because we are doing the opposite by adding 32!

OpenStudy (ja1):

Ok so the last equation had: \[(F - 32)\]so what is the opposite of that?

OpenStudy (ja1):

PERFECT

OpenStudy (ja1):

:D

OpenStudy (anonymous):

(F + 32)

OpenStudy (ja1):

See you got this just takes some practice

OpenStudy (anonymous):

LOL, it's just I need someone to help me through it.

OpenStudy (ja1):

Not quite, remember we are looking for F so it is (C + 32)

OpenStudy (ja1):

Yupp you just need a good understanding :)

OpenStudy (anonymous):

Thank you so much you're really helping me. So it's D.

OpenStudy (ja1):

No problem :) and medal would be appreciated ^,^

OpenStudy (ja1):

tu :)

OpenStudy (radar):